OpenStudy (anonymous):

what is 10 + 10 X 10 + ( 10 + ( 10 + 10 ) )

OpenStudy (anonymous):

140

OpenStudy (saifoo.khan):

140

OpenStudy (saifoo.khan):

+10xp.

OpenStudy (anonymous):

140

myininaya (myininaya):

20(40)=800

OpenStudy (anonymous):

10+10x10+(10+(20)) =10+10x10+(30) =10+100+30 =140

myininaya (myininaya):

where is the first open parenthesis?

OpenStudy (anonymous):

HA EVERYONE IS WRONGGGG IT WAS A TRICK! DO IT ON A CALCULATOR ITS 1200!

myininaya (myininaya):

(10+10)(10+(10+(10+10)) 20(10+(10+20)) 20(10+30) 20(40) 800 i get 800

OpenStudy (anonymous):

my calculator gives me 230 lol

OpenStudy (anonymous):

well you guys all suck at math, obviously. 1200 is the only correct answer.

myininaya (myininaya):

wait is that a X in between does two 10's i thought it was )( lol

myininaya (myininaya):

10+10*10+(10+(10+10)) 10+100+(10+20) 10+100+30 140 its 140

myininaya (myininaya):

1200 is extra wrong lol

OpenStudy (anonymous):

i cant tell if he's joking or not >.<

myininaya (myininaya):

me neither

myininaya (myininaya):

i see no way he got 1200

myininaya (myininaya):

i think he doesn't know because he said do it on a calculator lol

myininaya (myininaya):

yes my calculator agrees with 140 lol

OpenStudy (anonymous):

i typed it into a pocket cheapy calculator in order and it gave me 230 lol

myininaya (myininaya):

lol

OpenStudy (anonymous):

massetri... 10 + 10 X 10 + ( 10 + ( 10 + 10 ) ) Use BIDMAS Inner Brackets = 20 add outer brackets = 30 Multiplication 10 x 10 = 100 Add contents of brackets = 130 Add final 10 = 140 The logic in your calculator is spurious I believe!